OpenStudy (anonymous):

21 + 2a^4 - 4b^2 when a = -2 and b=3

OpenStudy (mathmale):

Have you included ALL of the instructions that came with this problem?

OpenStudy (anonymous):

yep

OpenStudy (mathmale):

I believe the instructions ask that you "Evaluate 21 + 2a^4 - 4b^2 when a=-2 and b=3." What does that mean to you?

OpenStudy (anonymous):

solve

OpenStudy (unklerhaukus):

\[21 + 2a^4 - 4b^2,\qquad a=-2,\qquad b=3\\ ~\\ =21+2(-2)^4-4(3)^2\\=\]

OpenStudy (anonymous):

No!

OpenStudy (anonymous):

I dont know really

OpenStudy (anonymous):

the answer is 17

OpenStudy (anonymous):

Can you explain how you got the answer cause I'm a little confused

OpenStudy (anonymous):

\[21 + 2*a^4 - 4*b^2\] \[=21 + 2*(-2)^4 - 4*(3)^2\] \[=21 + 2*16 - 4*9\] \[=21 + 32 - 36\] \[= 17\]

OpenStudy (anonymous):

Ohhh i see now

OpenStudy (anonymous):

good!

OpenStudy (anonymous):

Thanks

OpenStudy (anonymous):

you are welcome :)